OpenStudy (anonymous):

solve the quadratic equation by using the definition of square root. and write the solution in radical form.

OpenStudy (anonymous):

\[y ^{2}-80=0\]

OpenStudy (anonymous):

Add 80 to both sides to get y^2=80 Square root both sides to isolate a single y as y=sqrt(80) It technically has two answers. y=sqrt(80) and y=-sqrt(80)

OpenStudy (anonymous):

ohh ok, i wasn't sure if it was simplified to 40. ok thanks